OpenStudy (anonymous):

Find the point at which f(x)= x/(sqrt(9x-1) has a horizontal tangent line. x=?

OpenStudy (anonymous):

find 1º derivative and make it equal to 0. Solve that for x

OpenStudy (anonymous):

But because there are two x's do I have to use the qoutient theorem first?

OpenStudy (anonymous):

I came up with -1/2x(9x-1)^(-3/2)

OpenStudy (anonymous):

Then I came up with ((-9/2x)^2+(1/2x))^(-3/2)=0

OpenStudy (anonymous):

x=2/9 as dy/dx=0

OpenStudy (anonymous):

how did you find that?

OpenStudy (anonymous):

dy/dx=squareroot (9x-1)(1) -x*1*9/2*square root(9x-1) whole divided by (9x-1) =0 {as tangent horizontal dy/dx=o} c this is the quotient rule that we have applied and simplifying we get x=2/9

OpenStudy (anonymous):

if u like the result then let's chat i m feeling quite bored after long hours of studying....................................